# ExPLoRA: Parameter-Efficient Extended Pre-Training to Adapt Vision Transformers under Domain Shifts

### v3

@@ 1 Introduction
Pre-training foundation models (Bommasani et al., 2021) for natural language (Brown et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023; Jiang et al., 2024) and natural images (Oquab et al., 2023; He et al., 2022; Zhou et al., 2021; Rombach et al., 2022) has historically been computationally intensive, often limited to organizations with substantial resources. However, recent advancements in parameter-efficient fine-tuning (PEFT) techniques including low-rank adaptation (LoRA) and others (Hu et al., 2021; Zhang et al., 2023b; Chavan et al., 2023; Qiu et al., 2023; Liu et al., 2023; Jia et al., 2022) have sparked significant interest. These methods aim to adapt foundation models to downstream supervised-learning tasks using a small fraction (0.1%-10%) of the model’s trainable weights, with many based on the hypothesis that the required weight updates to the pre-trained model have a “low intrinsic rank" (Hu et al., 2021; Li et al., 2018; Aghajanyan et al., 2020).
@@
In this paper, we focus on visual foundation models (VFMs) such as DinoV2 or MAE (Oquab et al., 2023; He et al., 2022), which were trained on large-scale natural-image datasets. Despite the large investments in developing such models for natural images, they underperform when applied to other domains with visual data (e.g. medical or satellite images). For example, fine-tuning a model pre-trained on natural images on satellite image classification tasks is not as effective as fine-tuning one that was pre-trained on satellite images (Cong et al., 2022; Ayush et al., 2021). To bridge this gap, prevailing approaches invest similarly large levels of compute to pre-train VFMs on new domains, inspired by techniques developed for natural images (Cong et al., 2022; Reed et al., 2023; Khanna et al., 2024; Zhou et al., 2023; Moutakanni et al., 2024; Man et al., 2023).
@@
In this work, we challenge this paradigm (fig. 1), asking whether pre-training from scratch on each new domain is strictly necessary, since doing so is expensive (in compute and time) and precludes knowledge transfer from natural images. Instead, we wish to more efficiently leverage the rich semantic information encoded in natural-image VFMs to adapt them to new domains. Our proposed solution addresses these concerns using parameter-efficient self-supervised learning methods for domain adaptation.
@@
We introduce ExPLoRA, which generalizes vision foundation models to new domains by extending the pre-training phase with parameter-efficient techniques. We initialize a vision transformer (ViT) (Dosovitskiy et al., 2021) with weights pre-trained from natural-image datasets such as MAE or DinoV2. Selectively unfreezing 1-2 transformer blocks, we tune remaining weights with LoRA and continue unsupervised pre-training on the new domain. Subsequently fine-tuning with linear probing or LoRA on this new domain for supervised learning outperforms prior state-of-the-art (SoTA) approaches while training under 5-10% of the original weights. On satellite imagery, for example, we demonstrate an 8% improvement in linear probing top-1 accuracy, and even an improvement over prior SoTA fully pre-trained and fine-tuned techniques that required up to 16x more trainable parameters and 8x more compute (GPU hours). We conduct an extensive study on RGB, temporal, and multi-spectral satellite images, either matching or outperforming prior methods that fully pre-train from scratch. ExPLoRA also outperforms baselines and prior work on different domains such as wildlife, medical, and agricultural imagery on the WILDS (Koh et al., 2021) benchmark. Our contributions include:
@@
    Demonstrating ExPLoRA’s efficacy via ablation studies and by analyzing the differences in local (positional) and global (class) information encoded in the patch representations output by each ViT block (section 7).
@@ 2 Related Work
VFMs such as DinoV2 or masked autoencoders (MAE) that pre-train with self-supervised learning (SSL) have demonstrated remarkable performance across downstream tasks such as classification or semantic segmentation (Oquab et al., 2023; He et al., 2022; Grill et al., 2020; Chen et al., 2020). However, there has also been a rise in domain-specific VFMs (Cong et al., 2022; Reed et al., 2023; Moutakanni et al., 2024; Ma et al., 2024; Zhang et al., 2023a). For instance, SatMAE handles temporal or multi-spectral satellite image inputs. Since these models contain hundreds of millions of parameters, efficient adaptation to downstream tasks has become a key research focus.
@@
PEFT methods have gained widespread adoption for efficiently adapting large models by updating only a fraction of parameters, mitigating the prohibitive costs of full model tuning. LoRA learns low-rank weight updates to frozen weights, while other methods modify the frequency or number of trainable parameters per layer (Hu et al., 2021; Zhang et al., 2023b; Chavan et al., 2023; Pu et al., 2023). Others use multiplicative orthogonal updates (Qiu et al., 2023; Liu et al., 2023) or inject adapter modules (Steitz & Roth, 2024; Yin et al., 2023; Chen et al., 2022; Yin et al., 2024; Lian et al., 2022), effectively retaining pre-training knowledge in frozen weights. Visual prompt tuning (VPT) methods concatenate learnable prompt tokens to image patch sequences, trading improved fine-tuning performance with increased inference costs (Jia et al., 2022; Yoo et al., 2023; Pei et al., 2024; Han et al., 2023; Nie et al., 2023). ExPLoRA aims to supplement rather than replace these methods, and thus can be configured with any existing or future PEFT method for ViT fine-tuning.
@@
Domain adaptation enables models trained on a source domain to perform well on a different but related target domain. Traditional transformer-based methods address this via domain alignment, discriminative feature learning, cross-attention with pseudo-labels (Sun et al., 2022; Chuan-Xian et al., 2022; Zhu et al., 2023), or adversarial learning with self-refinement (Yang et al., 2023; Xu et al., 2021), typically requiring either labeled target data or source domain labels. Recent work explores adapting ViTs through different means: e.g., continual pre-training via masked image modeling (Mendieta et al., 2023) and scaled LoRA adapters (Scheibenreif et al., 2024) for satellite imagery. ExPLoRA builds on this direction, enabling parameter-efficient SSL directly on the target domain.
@@ 3 Background
DinoV2 (Oquab et al., 2023) is a robust SSL method for ViTs. Unlike MAE, DinoV2 features $f_{\theta}({\mathbf{x}})$ have demonstrated strong zero-shot performance, enabling adaptation to downstream tasks even with a frozen ViT backbone. During pre-training, DinoV2 maintains two copies of a ViT encoder: the student $f_{\theta}$ (trainable) and the teacher $f_{\theta^{\prime}}$, where $\theta^{\prime}$ is updated using an exponential-moving average of the student’s parameters $\theta$. The training objective incorporates a global, image-level loss from Dino (Caron et al., 2021), a patch-based loss from iBOT (Zhou et al., 2021), and regularizers including KoLeo (Delattre & Fournier, 2017) and Sinkhorn-Knopp centering (Caron et al., 2020).
@@
Low-rank adaptation (LoRA) (Hu et al., 2021) assumes that the weight update to change a set of unsupervised pre-trained weights to supervised fine-tuned weights lives in a low-rank subspace,
@@ 5 Method
where $f_{\theta}(\cdot;W_{S})$ is a ViT feature encoder parameterized by $\theta$ with initialization $W_{S}$, $g_{\psi}$ is a learnable decoder parameterized by $\psi$ (e.g., MAE/DinoV2 decoder), ${\mathcal{C}}_{S}$ is the unsupervised loss that was used for $W_{S}$ (e.g., reconstruction loss), and $\Theta({\mathcal{U}},r)$ restricts the parameter space to full-rank updates for blocks in ${\mathcal{U}}$ and LoRA rank-$r$ updates in ${\mathcal{L}}\setminus{\mathcal{U}}$. As we show in section 6, ExPLoRA can even outperform full pre-training on new domains from scratch.
@@
For DinoV2, we initialize a ViT-L with $W_{S}$ from DinoV2’s ViT encoder, without registers (Darcet et al., 2023). Since DinoV2 pre-trained checkpoints don’t include Dino or iBOT linear heads, we initialize a shared head $g_{\psi}$ from scratch, which is fully trained during extended pre-training.
@@
After running ExPLoRA, we receive a new unsupervised model $W_{T}^{\ast}=W_{S}+\Delta_{T}$ for the target domain $T$. $g_{\psi}$ (e.g., the Dino linear head or the MAE decoder) is discarded since it is not part of the ViT encoder $f$. Post-ExPLoRA, only $\Delta_{T}$, consisting of 1-2 unfrozen ViT blocks, LoRA matrices, and layer-normalization weights, are stored for each $T$– all of which can be merged into the original ViT, thus preserving architecture. Like LoRA, ExPLoRA significantly reduces additional storage requirements compared to fully training $W_{T}$ from scratch.
We can treat $W_{T}^{\ast}$ as any other fully pre-trained ViT $W_{T}$. Finally, solving for $\Delta^{(\tau)}$ on each labeled dataset ${\mathcal{Y}}^{(\tau)}_{T}$:
@@
where the ViT $f_{\theta}$ is now initialized with $W_{T}^{\ast}$, $h^{\tau}_{\phi}(\cdot)$ is a task-specific ViT decoding head with parameters $\phi$, and ${\mathcal{C}}^{\prime}$ represents a supervised-learning cost function.
For any PEFT method, we restrict $\theta\in\Theta(r)$ where $r$ controls the fraction of trainable weights in ${\mathcal{L}}$. (e.g., LoRA only on attention $Q,V$ matrices). For linear probing, $\Theta(0)=\emptyset$, so $(h^{\tau}_{\phi}\circ f_{\theta})({\mathbf{x}};W_{T}^{\ast})=\phi^{\top}f({\mathbf{x}};W_{T}^{\ast})$. For general fine-tuning, we optimize both $\theta\in\Theta(\infty)$ and $\phi$, modifying $h^{\tau}$ as per the task (e.g., a decoder head for segmentation). With $\Delta^{(\tau)}$, eq. 4 gives us our final model weights $W^{(\tau)}_{T}$, which can be used for classification, segmentation, detection etc.
@@ 6 Experiments
Our experimental results consist of a case study on satellite imagery (section 6.1), with an ablation study in section 6.1.2. We evaluate on multiple downstream tasks in sections 6.1.3, 6.1.4 and 6.2. Additional experiments and ablations are provided in appendix B and training hyperparameter and compute configurations are mentioned in appendix C. Our results achieve a new SoTA top 1 accuracy of 79.3% ($\uparrow$1.5%) on the competitive fMoW-RGB benchmark, outperforming fully pre-trained and fine-tuned models while using 6% of the ViT encoder parameters. We also achieve a $\uparrow$8.2improvement in linear probing accuracy on the same dataset. Across other satellite datasets, we match fully-pretrained prior state-of-the-art methods that required up to 8x more compute and 16x more trainable parameters and demonstrate competitive performance on WiLDS benchmark datasets as well.
@@ 6.1 Case Study: Satellite Imagery
We examine satellite images given their importance towards societal applications (Broader Impact) and since they represent a significant domain shift from natural images. There is a large and growing body of research on developing foundation models for satellite imagery from scratch (Cong et al., 2022; Reed et al., 2023; Tang et al., 2024), thus presenting a good benchmark for ExPLoRA.
@@ 6.1.1 RGB Satellite Images
Model PEFT
Pre-train/Fine-tune
#Params
Top 1
Acc.
ScaleMAE [54] Full 303.3M / 303.3M 77.80
SatMAE [17] Full 303.3M / 303.3M 77.78
SatMAE [17] LoRA-$r8$ [28] 303.3M / 0.8M 76.10
ScaleMAE [54] LoRA-$r8$ [28] 303.3M / 0.8M 78.01
GFM [44] LoRA-$r8$ [28] 303.3M / 0.8M 73.03
GDA [56] GDA-$r16$ [56] 8.5M / 8.5M 71.88
MAE [26] LoRA-$r8$ [28] - / 0.8M 76.21
M-[L]-$r64$ LoRA-$r8$ [28] 18.7M / 0.8M 76.55
DinoV2 [47] LoRA-$r8$ [28] - / 0.8M 78.08
DinoV2 [47] BOFT-$b2m8$ [40] - / 0.9M 72.40
DinoV2 [47] Mona [69] - / 7.1M 72.80
DinoV2 [47] VPT-100 [29] - / 0.4M 77.29
DinoV2 [47] GVPT-100 [70] - / 0.4M 76.22
DinoV2 [47] AdaLoRA-$r8$ [72] - / 1.2M 78.87
DinoV2 [47] Adapter+ [59] - / 1.4M 78.16
DinoV2 [47] SA2VP [49] - / 1.1M 77.53
D-[L]-$r64$ SA2VP [49] 18.7M / 1.1M 78.51
D-[L]-$r64$ LoRA-$r8$ [28] 18.7M / 0.8M 79.28
Table 1: Results on the fMoW-RGB validation dataset. The “Pre-train / Fine-tune #Params" refer to the trainable parameters of the ViT-L encoder required on the new domain, i.e. satellite images. M-[L]-$r64$ and D-[L]-$r64$ refer to ExPLoRA models initialized with MAE and DinoV2 weights, respectively (section 5).
@@
GASSL [2] ResNet 68.32
SatMAE [17] ViT-L 65.94
ScaleMAE [54] ViT-B 67.30
CrossScaleMAE [61] ViT-B 69.20
DinoV2 [47] ViT-L 67.60
DinoV2$\dagger$ [47] ViT-L 69.00
D-[L]-$r64$ ViT-L 76.86
D-[L]-$r64$$\dagger$ ViT-L 77.48
@@
We compare our results in table 1 against both prior fully pre-trained SoTA foundation models as well as PEFT techniques applied on ViTs pre-trained with MAE and/or DinoV2 weights. Our results demonstrate that D-ExPLoRA-$\tt{[L]}$-$r64$ is SoTA in terms of fMoW-RGB average accuracy at 79.28%. ExPLoRA outperforms techniques that require fully and/or continually pre-training ViTs on fMoW while using 6% of the original ViT encoder parameters. Further experiments with MAE are in B.6.
ExPLoRA-initializations with LoRA fine-tuning outperform other unsupervised initializations paired with PEFT techniques by 1-3%, including SoTA matrix-adaptation methods like AdaLoRA (Zhang et al., 2023b), BOFT (Liu et al., 2023), VPT approaches such as GVPT (Yoo et al., 2023), SA2VP (Pei et al., 2024), and adapter methods like Adapter+ (Steitz & Roth, 2024). We also outperform satellite image continual pre-training methods such as GFM (Mendieta et al., 2023) and GDA (Scheibenreif et al., 2024) by 6%. Additionally, applying SA2VP to ExPLoRA-initialized ViTs improves performance over DinoV2 by  1%, showcasing ExPLoRA’s compatibility with other PEFT methods and its versatility as an initialization for new domains.

### v4

@@ 1 Introduction
Pre-training foundation models (Bommasani et al., 2021) for natural language (Brown et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023) and natural images (He et al., 2022; Rombach et al., 2022; Oquab et al., 2024) has historically been computationally intensive, often limited to organizations with substantial resources. However, recent advancements in parameter-efficient fine-tuning (PEFT) techniques including low-rank adaptation (LoRA) and others (Hu et al., 2022; Jia et al., 2022; Qiu et al., 2023) have sparked significant interest. These methods aim to adapt foundation models to downstream supervised-learning tasks using a small fraction (0.1%-10%) of the model’s trainable weights, with many based on the hypothesis that the required weight updates to the pre-trained model have a “low intrinsic rank" (Li et al., 2018; Aghajanyan et al., 2020).
@@
In this paper, we focus on visual foundation models (VFMs) such as MAE or DinoV2 (He et al., 2022; Oquab et al., 2024) that were trained on a large amount of natural images. Despite the big investments in developing VFMs for natural images, they underperform when applied to other domains with visual data (e.g. medical or satellite images). For example, fine-tuning a model pre-trained on natural images on satellite image classification tasks is not as effective as fine-tuning one that was pre-trained on satellite images (Ayush et al., 2021; Cong et al., 2022). To bridge this gap, prevailing approaches invest similarly large levels of compute to pre-train VFMs on new domains, inspired by techniques developed for natural images (Cong et al., 2022; Zhou et al., 2023; Moutakanni et al., 2024; Khanna et al., 2024).
@@
In this work, we challenge this paradigm (fig. 1), asking whether pre-training from scratch on each new domain is strictly necessary, since doing so is expensive (in compute and time) and precludes knowledge transfer from natural images. Instead, we wish to more efficiently leverage the rich semantic information encoded in natural-image VFMs to adapt them to new domains. Our proposed solution addresses these concerns using parameter-efficient self-supervised learning methods for transfer learning.
@@
We introduce ExPLoRA, which generalizes vision foundation models to new domains by extending the pre-training phase with parameter-efficient techniques. We initialize a vision transformer (ViT) (Dosovitskiy et al., 2021) with weights pre-trained from natural-image datasets such as MAE or DinoV2. Selectively unfreezing 1-2 transformer blocks, we tune remaining weights with LoRA and continue unsupervised pre-training on the new domain. Subsequently fine-tuning with linear probing or LoRA on this new domain for supervised learning outperforms prior state-of-the-art (SoTA) approaches while training under 5-10% of the original weights. On satellite imagery, we demonstrate an 8% improvement in linear probing top-1 accuracy, and even an improvement over prior SoTA fully pre-trained and fine-tuned techniques that required up to 16x more trainable parameters and 8x more compute (GPU hours). We conduct an extensive study on RGB, temporal, and multi-spectral satellite images, either matching or outperforming prior methods that fully pre-train from scratch. ExPLoRA also outperforms previous work on different domains such as wildlife, medical, and agricultural imagery on the WILDS (Koh et al., 2021) benchmark. Our contributions include:
@@
    Demonstrating ExPLoRA’s efficacy via ablation studies and by analyzing improvements in local (positional) and global (class) information encoded in the patch representations output by each ViT block (section 7).
@@ 2 Related Work
VFMs such as DinoV2 or masked autoencoders (MAE) that pre-train with self-supervised learning (SSL) have demonstrated remarkable performance across downstream tasks such as classification or semantic segmentation (Grill et al., 2020; Chen et al., 2020; He et al., 2022; Oquab et al., 2024). However, there has also been a rise in domain-specific VFMs (Cong et al., 2022; Man et al., 2023; Zhang et al., 2023b; Ma et al., 2024; Moutakanni et al., 2024). For instance, SatMAE handles temporal or multi-spectral satellite image inputs. Since these models contain hundreds of millions of parameters, efficient adaptation to downstream tasks has become a key research focus.
@@
PEFT methods have gained widespread adoption for efficiently adapting large models by updating only a fraction of parameters, mitigating the prohibitive costs of full model tuning. LoRA learns low-rank weight updates to frozen weights, while other methods modify the frequency or number of trainable parameters per layer (Hu et al., 2022; Zhang et al., 2023c; Chavan et al., 2023; Pu et al., 2023). Others use multiplicative orthogonal updates (Qiu et al., 2023; Liu et al., 2024) or inject adapter modules (Chen et al., 2022; Lian et al., 2022; Yin et al., 2023, 2024; Steitz & Roth, 2024), effectively retaining pre-training knowledge in frozen weights. Visual prompt tuning (VPT) methods concatenate learnable prompt tokens to image patch sequences, trading improved fine-tuning performance with increased inference costs (Jia et al., 2022; Yoo et al., 2023; Han et al., 2023; Nie et al., 2023; Tsai et al., 2023; Pei et al., 2024). ExPLoRA aims to supplement rather than replace these methods, and thus can be configured with any existing or future PEFT method for ViT fine-tuning.
@@
Domain adaptation enables models trained on a source domain to perform well on a different but related target domain. Traditional transformer-based methods address this via domain alignment, discriminative feature learning, cross-attention with pseudo-labels (Sun et al., 2022; Chuan-Xian et al., 2022; Zhu et al., 2023), or adversarial learning with self-refinement (Xu et al., 2021; Yang et al., 2023), typically requiring labeled target data or source domain labels. Test-time adaptation methods (Wang et al., 2021; Gao et al., 2022; Zhang et al., 2023a) adapt models without target domain labels, but assume a shared label space between domains and thus require models trained with supervised learning on the source domain. While  Reed et al. (2022) showed that sequential self-supervised pre-training on source and target datasets improves convergence for supervised tasks, their study was limited to ResNet-50 with MoCo and required full model training. Recent work adapts ViTs through different means: e.g., continual pre-training via masked image modeling (Mendieta et al., 2023) and scaled LoRA adapters (Scheibenreif et al., 2024) for satellite imagery. ExPLoRA builds on this direction, enabling parameter-efficient SSL directly on the target domain.
@@ 3 Background
DinoV2 (Oquab et al., 2024) is a robust SSL method for ViTs. Unlike MAE, DinoV2 features $f_{\theta}({\mathbf{x}})$ have demonstrated strong zero-shot performance, enabling adaptation to downstream tasks even with a frozen ViT backbone. During pre-training, DinoV2 maintains two copies of a ViT encoder: the student $f_{\theta}$ (trainable) and the teacher $f_{\theta^{\prime}}$, where $\theta^{\prime}$ is updated using an exponential-moving average of the student’s parameters $\theta$. The training objective incorporates a global, image-level loss from Dino (Caron et al., 2021), a patch-based loss from iBOT (Zhou et al., 2021), and regularizers including KoLeo (Delattre & Fournier, 2017) and Sinkhorn-Knopp centering (Caron et al., 2020).
@@
Low-rank adaptation (LoRA) (Hu et al., 2022) assumes that the weight update to change a set of unsupervised pre-trained weights to supervised fine-tuned weights lives in a low-rank subspace,
@@ 5 Method
where $f_{\theta}(\cdot;W_{S})$ is a ViT feature encoder parameterized by $\theta$ with initialization $W_{S}$, $g_{\psi}$ is a learnable decoder parameterized by $\psi$ (e.g., MAE/DinoV2 decoder), ${\mathcal{C}}_{S}$ is the unsupervised loss that was used for $W_{S}$ (e.g., reconstruction loss), and $\Theta({\mathcal{U}},r)$ restricts the trainable parameter space to full-rank updates for blocks in ${\mathcal{U}}$ and LoRA rank-$r$ updates in ${\mathcal{L}}\setminus{\mathcal{U}}$. Similar to Goyal et al. (2023), we find that using the same unsupervised loss ${\mathcal{C}}_{S}$ as was used for $W_{S}$ is beneficial. As we show in section 6, ExPLoRA can even outperform full pre-training on new domains from scratch.
@@
For DinoV2, we initialize a ViT-L with $W_{S}$ from DinoV2’s ViT encoder, without registers (Darcet et al., 2024). Since DinoV2 pre-trained checkpoints don’t include Dino or iBOT linear heads, we initialize a shared head $g_{\psi}$ from scratch, which is fully trained during extended pre-training.
@@
After running ExPLoRA, we receive a new unsupervised model $W_{T}^{\ast}=W_{S}+\Delta_{T}$ for the target domain $T$ that functions as any other pre-trained ViT $W_{T}$. $g_{\psi}$ (e.g., the Dino/MAE decoder) is discarded as it is not part of the ViT encoder $f$. Only $\Delta_{T}$, consisting of 1-2 unfrozen ViT blocks, LoRA matrices, and layer-normalization weights, are stored for each $T$– all of which can be merged into the original ViT, thus preserving architecture. Like LoRA, ExPLoRA significantly reduces additional storage requirements compared to fully pre-training $W_{T}$.
Finally, solving for $\Delta^{(\tau)}$ on each labeled dataset ${\mathcal{Y}}^{(\tau)}_{T}$:
@@
where the ViT $f_{\theta}$ is now initialized with $W_{T}^{\ast}$, $h^{(\tau)}_{\phi}(\cdot)$ is a task-specific ViT decoding head with parameters $\phi$, and ${\mathcal{C}}^{\prime}$ represents a supervised-learning loss function.
For any PEFT method, we allow $\theta\in\Theta(r)$, where $r$ restricts the trainable parameter space $\Theta$ of ${\mathcal{L}}$ (e.g., LoRA only on attention $Q,V$ matrices). For linear probing, $\Theta(0)=\emptyset$, so $(h^{(\tau)}_{\phi}\circ f_{\theta})({\mathbf{x}};W_{T}^{\ast})=\phi^{\top}f({\mathbf{x}};W_{T}^{\ast})$. For general fine-tuning, we optimize both $\theta\in\Theta$ and $\phi$ unrestricted, modifying $h^{(\tau)}$ as per the task (e.g., a decoder head for segmentation). With $\Delta^{(\tau)}$, eq. 4 gives us our final model weights $W^{(\tau)}_{T}$, which can be used for classification, segmentation, detection etc.
@@ 6 Experiments
Our experimental results consist of a case study on satellite imagery (section 6.1), with an ablation study in section 6.1.2. We evaluate on multiple downstream tasks in sections 6.1.3, 6.1.4 and 6.2. Additional experiments and ablations are provided in appendix B and training hyperparameter and compute configurations are mentioned in appendix C.
We report both performance metrics and computational requirements for our experiments. As a pre-training technique, ExPLoRA offers an efficient alternative to full domain-specific pre-training. Thus, unlike task-specific fine-tuning, its compute costs are amortized across all downstream applications of the resulting model including feature extraction, linear probing, and various supervised tasks.
Our results achieve a new SoTA top 1 accuracy of 79.3% ($\uparrow$1.5%) on the competitive fMoW-RGB benchmark, outperforming fully pre-trained and fine-tuned models while using 6% of the ViT encoder parameters and requiring only 100 GPU hours compared to 960+ hours for full pre-training. We also achieve a $\uparrow$8.2improvement in linear probing accuracy on the same dataset. Across other satellite datasets, ExPLoRA matches or exceeds fully-pretrained prior state-of-the-art methods while requiring 8x-10x less compute and 16x fewer trainable parameters and demonstrates competitive performance on WiLDS benchmark datasets as well.
@@ 6.1 Case Study: Satellite Imagery
We examine satellite images given their importance towards societal applications (Impact Statement) and since they represent a significant domain shift from natural images. There is a large and growing body of research on developing foundation models for satellite imagery from scratch (Cong et al., 2022; Reed et al., 2023; Tang et al., 2024), thus presenting a good benchmark for ExPLoRA.
Model PEFT
Pre-Train
Params
Fine-Tune
Params
Pre-Train
GPU hours
Fine-Tune
GPU hours
Top 1 Acc.
1          ScaleMAE [56] Full 303.3M 303.3M 960 610 77.80
2          SatMAE [17] Full 303.3M 303.3M 960 610 77.78
3          SatMAE [17] LoRA-$r8$ [30] 303.3M 0.8M 960 220 76.10
4          ScaleMAE [56] LoRA-$r8$ [30] 303.3M 0.8M 960 220 78.01
5          GFM [45] LoRA-$r8$ [30] 303.3M 0.8M 900 220 73.03
6          GDA [58] GDA-$r16$ [58] 8.5M 8.5M 310 310 71.88
7          MAE [28] LoRA-$r8$ [30] 0.8M 220 76.21
8          M-[L]-$r64$ LoRA-$r8$ [30] 18.7M 0.8M 80 220 76.55
9          DinoV2 [48] LoRA-$r8$ [30] 0.8M 220 78.08
10          DinoV2 [48] BOFT-$b2m8$ [41] 0.9M 230 72.40
11          DinoV2 [48] Mona [73] 7.1M - 220 72.80
12          DinoV2 [48] VPT-100 [31] 0.4M 500 77.29
13          DinoV2 [48] GVPT-100 [74] 0.4M 500 76.22
14          DinoV2 [48] AdaLoRA-$r8$ [77] 1.2M 220 78.87
15          DinoV2 [48] Adapter+ [61] 1.4M 220 78.16
16          DinoV2 [48] SA2VP [50] 1.1M 410 77.53
17          D-[L]-$r64$ SA2VP [50] 18.7M 1.1M 100 410 78.51
18          D-[L]-$r64$ LoRA-$r8$ [30] 18.7M 0.8M 100 220 79.28
Table 1: Results on the fMoW-RGB validation dataset. "Pre-train / Fine-tune Params" refer to trainable parameters of the ViT-L encoder required on the new domain, i.e. satellite images. M-[L]-$r64$ and D-[L]-$r64$ refer to ExPLoRA models initialized with MAE and DinoV2 weights, respectively (section 5). Our measurements for GPU hours use standardized hardware platforms for fair comparison.
@@ 6.1.1 RGB Satellite Images
1         GASSL [2] ResNet 68.32
2         SatMAE [17] ViT-L 65.94
3         ScaleMAE [56] ViT-B 67.30
4         CrossScaleMAE [63] ViT-B 69.20
5         DinoV2 [48] ViT-L 67.60
6         DinoV2$\dagger$ [48] ViT-L 69.00
7         D-[L]-$r64$ ViT-L 76.86
8         D-[L]-$r64$$\dagger$ ViT-L 77.48
@@
We compare our results in table 1 against both prior fully pre-trained SoTA foundation models as well as PEFT techniques applied on ViTs pre-trained with MAE and/or DinoV2 weights. Our results demonstrate that D-ExPLoRA-$\tt{[L]}$-$r64$ is SoTA in terms of fMoW-RGB average accuracy at 79.28%. ExPLoRA outperforms techniques that require fully and/or continually pre-training ViTs on fMoW while using 6% of the original ViT encoder parameters and 8x less compute. Further experiments with MAE are in B.6.
ExPLoRA-initializations with LoRA fine-tuning outperform other unsupervised initializations paired with PEFT techniques by 1-3%, including SoTA matrix-adaptation methods like AdaLoRA (Zhang et al., 2023c), BOFT (Liu et al., 2024), VPT approaches such as GVPT (Yoo et al., 2023), SA2VP (Pei et al., 2024), and adapter methods like Adapter+ (Steitz & Roth, 2024). We also outperform satellite image continual pre-training methods such as GFM (Mendieta et al., 2023) and GDA (Scheibenreif et al., 2024) by 6%. Additionally, applying SA2VP to ExPLoRA-initialized ViTs improves performance over DinoV2 by  1% (rows 16 vs 17), showcasing ExPLoRA’s compatibility with other PEFT methods and its versatility as an initialization for new domains.